OpenStudy (anonymous):

which is the sum of the solutions, x + y, for the systems of linear equations 7x + 3y = -12 2x + 5y = 38

OpenStudy (anonymous):

go help that guy with the ellipse!

OpenStudy (anonymous):

after i complete the square conics make me dizzy

OpenStudy (anonymous):

wut?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

lets do this mulitiply the top equation by -5, bottom by 3 and see what we get

OpenStudy (anonymous):

\[-35x-15y=60\] \[6x+15y=114\]

OpenStudy (anonymous):

check my arithmetic but i think it is right. now add to get \[-29x=174\]

OpenStudy (anonymous):

making \[x=\frac{174}{-29}=-6\]

OpenStudy (anonymous):

if x = -6 then the first equation give \[-42+3y=-12\] \[3y=30\] \[y=10\]

OpenStudy (anonymous):

wait slow down down after i get -29x = 174, where do i go from there and remember im not good at math so use easy words

OpenStudy (anonymous):

so the sum is 10-6=4

OpenStudy (anonymous):

was it clear how to get -29x=174? because that is all the work

OpenStudy (anonymous):

yah but aftr that im confused

OpenStudy (anonymous):

ok that was the hard part. the easy part is this: if \[-29x=174\] then this is a one step equation to solve. to get x by itself divide both sides by -29

OpenStudy (anonymous):

the -29 will cancel on the left, leaving x by itself, and on the right you get whatever 174/-29 is

OpenStudy (anonymous):

i used a calculator

OpenStudy (anonymous):

was that part ok?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so now we know that x = -6 because that is what i got from 174/-29

OpenStudy (anonymous):

now i take the -6 and substitute it in to either of your first two equations for x

OpenStudy (anonymous):

you can use either equation. i chose the first one. where i saw an x i put a -6

OpenStudy (anonymous):

it was \[7x + 3y = -12\]

OpenStudy (anonymous):

so i replace x by -6 and get \[7\times -6+3y=-12\]

OpenStudy (anonymous):

ok so if the final ans is 4 then i know how to to do it

OpenStudy (anonymous):

i am not done yet. i need x which is -6. then i need y so i have to solve the above equation for y

OpenStudy (anonymous):

\[-42+3y=-12\] \[3y=-12+42\] \[3y=30\] \[y=\frac{30}{3}\] \[y=10\]

myininaya (myininaya):

right! x+y=-6+10=4 gj satellite

OpenStudy (anonymous):

now i know x = -6 and i know y = 10 so i know that x+y is -6+10=4

OpenStudy (anonymous):

yay i got it